OpenStudy (anonymous):

Find the Value of COS 11pi/6 Show work So i can Understand how to do this . Thankssssss :)

OpenStudy (anonymous):

Cos is the x axis and Sin is the y axis. If you look at a unit circle I attached, follow it around until you reach 11pi/6 (circled in red), which corresponds to root3/2 (circled in green).

OpenStudy (anonymous):

So my answer is what you circled in green ?

OpenStudy (anonymous):

Yes.\[\frac{\left(\frac{11\pi }{6}\right)}{(2\pi )}360{}^{\circ}=330 {}^{\circ} \]or \[-30{}^{\circ} \]\[\text{Cos}[-30{}^{\circ}]=\frac{\sqrt{3}}{2} \]

OpenStudy (anonymous):

Yep.